# ARC Prize 2024: Technical Report

### v1

@@ Abstract
As of December 2024, the ARC-AGI benchmark is five years old and remains unbeaten. We believe it is currently the most important unsolved AI benchmark in the world because it seeks to measure generalization on novel tasks – the essence of intelligence – as opposed to skill at tasks that can be prepared for in advance. This year, we launched ARC Prize, a global competition to inspire new ideas and drive open progress towards AGI by reaching a target benchmark score of 85%. As a result, the state-of-the-art score on the ARC-AGI private evaluation set increased from 33% to 55.5%, propelled by several frontier AGI reasoning techniques including deep learning-guided program synthesis and test-time training. In this paper, we survey top approaches, review new open-source implementations, discuss the limitations of the ARC-AGI-1 dataset, and share key insights gained from the competition.
@@ 2.3 Paper Awards
    • Franzen et al., (the ARChitects): “The LLM ARChitect: Solving the ARC Challenge Is a Matter of Perspective”
@@
While we have a long way to go to reach AGI, we’re excited that ARC Prize has catalyzed several new open source frontier AGI reasoning approaches, in particular test-time training, an approach we first observed in use by Jack Cole in 2023 and subsequently popularized this year by ARC Prize.
@@ 3.2 Test-Time Training
The classical deep learning paradigm, exemplified by LLMs from the 2022-2023 period, involves first training a model on a large dataset and then performing inference with a frozen version of the model. However, solving ARC-AGI requires going beyond simple fetching and application of memorized patterns – it necessitates the ability to adapt to the specific task at hand at test time. This has led to the emergence of test-time training (TTT), also known as test-time fine-tuning (TTFT), as a dominant approach in LLM-based solutions for ARC-AGI. Today, all top LLM-based transduction approaches for ARC-AGI leverage TTT, and there does not exist any static inference-style transduction solution that scores above 10%. This stark gap highlights the inability of the classical deep learning paradigm to generalize to novel tasks.
@@
  • Data augmentation and alternative datasets: Given the limited size of the ARC-AGI-1 dataset, successful TTT relies heavily on increasing the amount of ARC-AGI-like data available for pretraining, either by using alternative datasets like BARC (a dataset of 400,000 additional ARC-like tasks released by Ellis et al.  [19], Re-ARC (a programmatic implementation of the ARC-AGI training set allowing infinite sampling of new instances of the 400 training tasks, released by Hodel  [13]) or carefully designed data augmentations. The ARChitects  [10], for example, introduced novel augmentations and a selection criterion based on the stability of generated solutions under these augmentations.
@@ 3.3 Combining Program Synthesis with Transduction
As soon as transduction-based approaches started scoring above zero (in late 2023, pioneered by Jack Cole) researchers noticed that program search and transduction were able to solve significantly distinct sets of tasks. This issue was later investigated in detail by Li et al. in “Combining Induction and Transduction for Abstract Reasoning”  [19].
Today, all top scores (e.g., Akyürek and Berman on the public leaderboard, the ARChitects, Barbadillo, and MindsAI on the Kaggle leaderboard) use a combination of transduction and induction. The best transduction-only and induction-only single submissions score around 40%, so only an ensemble of both can compete for the state of the art.
@@ 6.1 The ARC-AGI Ecosystem
  • BARC - Tools for generating synthetic ARC-AGI problems: GitHub Repository
@@ 6.2 Acknowledgments
ARC Prize builds on the legacy of earlier ARC-AGI competitions: the 2020 competition on Kaggle as well as the 2022 and 2023 “ARCathons”, which were a collaboration between François Chollet and the Davos-based non-profit AI lab, Lab42. We are grateful to the Lab42 team – Rolf Pfister, Oliver Schmid, and Hansueli Jud – for their expertise and support in facilitating a smooth transition for the ARC-AGI community. Their contributions were significant in facilitating a bigger, bolder competition and advancing initiatives like ARC-AGI-2.

### v2

@@ Abstract
As of December 5, 2024, the ARC-AGI benchmark is five years old and remains unbeaten. We believe it is currently the most important unsolved AI benchmark in the world because it seeks to measure generalization on novel tasks – the essence of intelligence – as opposed to skill at tasks that can be prepared for in advance. This year, we launched ARC Prize, a global competition to inspire new ideas and drive open progress towards AGI by reaching a target benchmark score of 85%. As a result, the state-of-the-art score on the ARC-AGI private evaluation set increased from 33% to 55.5%, propelled by several frontier AGI reasoning techniques including deep learning-guided program synthesis and test-time training. In this paper, we survey top approaches, review new open-source implementations, discuss the limitations of the ARC-AGI-1 dataset, and share key insights gained from the competition.
@@ 2.3 Paper Awards
    • Franzen et al., (the ARChitects): “The LLM ARChitect: Solving ARC-AGI Is A Matter of Perspective”
@@
While we have a long way to go to reach AGI, we’re excited that ARC Prize has catalyzed several new open source frontier AGI reasoning approaches, in particular test-time training, an approach we first observed in use by Jack Cole and Mohamed Osman in 2023 and subsequently popularized this year by ARC Prize.
@@ 3.2 Test-Time Training
The classical deep learning paradigm, exemplified by LLMs from the 2022-2023 period, involves first training a model on a large dataset and then performing inference with a frozen version of the model. However, solving ARC-AGI requires going beyond simple fetching and application of memorized patterns – it necessitates the ability to adapt to the specific task at hand at test time. This has led to the emergence of test-time training (TTT), also known as test-time fine-tuning (TTFT), as a dominant approach in LLM-based solutions for ARC-AGI. Today, all top LLM-based transduction approaches for ARC-AGI leverage TTT, and there does not exist any static inference-style transduction solution that scores above 11%. This stark gap highlights the inability of the classical deep learning paradigm to generalize to novel tasks.
@@
  • Data augmentation and alternative datasets: Given the limited size of the ARC-AGI-1 dataset, successful TTT relies heavily on increasing the amount of ARC-AGI-like data available for pretraining, either by using alternative datasets like ARC-Heavy/ARC-Potpourri (a combined dataset comprising 400,000 additional ARC-like tasks released by Ellis et al.  [19]) and Re-ARC (a programmatic implementation of the ARC-AGI training set allowing infinite sampling of new instances of the 400 training tasks, released by Hodel  [13]), or by carefully designed data augmentations. The ARChitects  [10], for example, introduced novel augmentations and a selection criterion based on the stability of generated solutions under these augmentations.
@@ 3.3 Combining Program Synthesis with Transduction
As soon as transduction-based approaches started scoring above zero (in late 2023, pioneered by Jack Cole and Mohamed Osman) researchers noticed that program search and transduction were able to solve significantly distinct sets of tasks. This issue was later investigated in detail by Li et al. in “Combining Induction and Transduction for Abstract Reasoning”  [19].
Today, all top scores (e.g., Akyürek and Berman on the public leaderboard, and the ARChitects and Barbadillo on the Kaggle leaderboard) use a combination of transduction and induction. The best transduction-only and induction-only single submissions score around 40%, so only an ensemble of both can compete for the state of the art.
@@ 6.1 The ARC-AGI Ecosystem
  • Bootstrapping ARC - Tools for generating synthetic ARC-AGI problems: GitHub Repository
@@ 6.2 Acknowledgments
ARC Prize builds on the legacy of earlier ARC-AGI competitions: the 2020 competition on Kaggle as well as the 2022 and 2023 “ARCathons”, which were a collaboration between François Chollet and the Davos-based non-profit AI lab, Lab42. We are grateful to the Lab42 team – Pascal Kaufmann, Rolf Pfister, Oliver Schmid, and Hansueli Jud – for their expertise and support in facilitating a smooth transition for the ARC-AGI community. Their contributions were significant in facilitating a bigger, bolder competition and advancing initiatives like ARC-AGI-2.
@@
ARC Prize would have not been possible without the support of the Kaggle team. Kaggle plays a critical role in the Artificial and Machine Learning ecosystem. We would like to thank Addison Howard, Walter Reade, Maggie Demkin and Elizabeth Park for the support throughout the 2024 competition.